OpenStudy (anonymous):

I need help! the circumfrence of a spherical candle is 15.7 inches. What is the volume of wax in the candle?

OpenStudy (anonymous):

4pir=15.7 so r =15.7/4pi volume=4pir^3/3 use this value i=of r to find the volume :)

OpenStudy (amistre64):

lol....its a ball of wax

OpenStudy (anonymous):

I still am trying to figure out how to solve it!

OpenStudy (amistre64):

find the radius bu cutting it in half; and taking the radius of the circumference of the circle

OpenStudy (amistre64):

2pir = 15.7 r = 15.7/2pi

OpenStudy (amistre64):

umm.... 4pi r = surface area; not circumference if i recall correctly

OpenStudy (amistre64):

then the volume of the wax ball is: \[\frac{4\pi (\frac{15.7}{2\pi})^3}{3}\]

OpenStudy (anonymous):

i don't think there's anything called a cicurmference for a 3d object...i think he meant surface area...

OpenStudy (amistre64):

volume should be abt; 65.35....

OpenStudy (amistre64):

the circumference of a 3d ball is just the measure of the 'great cirlce'

OpenStudy (anonymous):

cubed or squared?

OpenStudy (amistre64):

volume is in cubed units

OpenStudy (amistre64):

inches^3

OpenStudy (anonymous):

Thank you guys for helping me :)

OpenStudy (anonymous):

awesome thanks

OpenStudy (amistre64):

youre welcome :)